# Regret Analysis of Average-Reward Unichain MDPs via an Actor-Critic Approach

### v1

@@ D.3 Proof of Lemma 3
D.4 Bound on the Hitting time of the Recurrent Class
@@
E Analysis of a General Linear Recursion
@@
$(A_{8})\quad x^{\top}Px\geq\lambda_{P}\|x\|^{2},\quad(A_{9})\quad\mathrm{Ker}(P)\subseteq\mathrm{Ker}(\widehat{P}_{h})$
for all $x\in\mathrm{Ker}(P)_{\perp}$ and $h\geq 1$. Let $x^{*}=P^{\dagger}q$, where $P^{\dagger}$ is the Moore-Penrose pseudoinverse of $P$ and $\Pi$ denote the projection operator onto the space $\mathrm{Ker}(P)_{\perp}$, i.e., the orthogonal complement of the kernel (null space) of $P$.
@@ E.1 Proof of Theorem 4(a)
Consider a matrix $A$ such that $x^{\top}Ax\geq\lambda\|x\|^{2}$ for all $x\in(\mathrm{Ker}(A))_{\perp}$. Then, we have
@@
Let $\Pi$ denote the projection operator onto the space $\mathrm{Ker}(P)_{\perp}$, i.e., the orthogonal complement of the kernel (null space) of $P$. Then the following properties hold:
@@
    If $\widetilde{P}$ is such that $\mathrm{Ker}(P)\subseteq\mathrm{Ker}(\widetilde{P})$, then $\widetilde{P}\Pi=\widetilde{P}$.
@@
For the second property, let $\Pi_{\perp}$ denote the projection operator onto $\mathrm{Ker}(P)$. By the orthogonal decomposition of the space, we have
$\Pi+\Pi_{\perp}=I,$
@@
$\widetilde{P}x=\widetilde{P}(\Pi x+\Pi_{\perp}x)=\widetilde{P}\Pi x,$
where the last equality holds since $\Pi_{\perp}x\in\mathrm{Ker}(P)\subseteq\mathrm{Ker}(\widetilde{P})$ and thus $\widetilde{P}\Pi_{\perp}x=0$. This establishes the result. ∎
@@
$\begin{split}&\operatorname{\mathbb{E}}\langle\Pi(x_{h}-x^{*})-\bar{\beta}\Pi P(x_{h}-x^{*}),\bar{\beta}\Pi M_{h+1}\rangle\\ &\quad=\operatorname{\mathbb{E}}\langle\Pi(x_{h}-x^{*})-\bar{\beta}\Pi P(x_{h}-x^{*}),\bar{\beta}\operatorname{\mathbb{E}}[\Pi M_{h+1}|x_{h}]\rangle\\ &\quad\overset{(a)}{\leq}\frac{\bar{\beta}\lambda_{P}}{2}\operatorname{\mathbb{E}}\|\Pi(x_{h}-x^{*})-\bar{\beta}\Pi P(x_{h}-x^{*})\|^{2}+\frac{2\bar{\beta}}{\lambda_{P}}\operatorname{\mathbb{E}}\|\Pi\operatorname{\mathbb{E}}[M_{h+1}|x_{h}]\|^{2}\\ \end{split}$ (60)
@@
$\begin{split}&\operatorname{\mathbb{E}}\|\Pi(x_{h}-x^{*})-\bar{\beta}\Pi P(x_{h}-x^{*})\|^{2}+\operatorname{\mathbb{E}}\langle\Pi(x_{h}-x^{*})-\bar{\beta}\Pi P(x_{h}-x^{*}),\bar{\beta}\Pi M_{h+1}\rangle\\ &\leq\left(1+\frac{\bar{\beta}\lambda_{P}}{2}\right)\operatorname{\mathbb{E}}\|\Pi(x_{h}-x^{*})-\bar{\beta}\Pi P(x_{h}-x^{*})\|^{2}+\frac{2\bar{\beta}}{\lambda_{P}}\|\Pi\operatorname{\mathbb{E}}[M_{h+1}|x_{h}]\|^{2}\\ &=\left(1+\frac{\bar{\beta}\lambda_{P}}{2}\right)\left(1-\lambda_{P}\beta+\beta^{2}\Lambda_{P}\right)\operatorname{\mathbb{E}}\|\Pi(x_{h}-x^{*})\|^{2}+\frac{2\bar{\beta}}{\lambda_{P}}\|\Pi\operatorname{\mathbb{E}}[M_{h+1}|x_{h}]\|^{2}\\ &=\left(1-\frac{\bar{\beta}\lambda_{P}}{2}+\frac{\bar{\beta}^{2}(2\Lambda_{P}-\lambda_{P}^{2})}{2}+\frac{\bar{\beta}^{3}\Lambda_{P}\lambda_{P}}{2}\right)\operatorname{\mathbb{E}}\|\Pi(x_{h}-x^{*})\|^{2}+\frac{2\bar{\beta}}{\lambda_{P}}\|\Pi\operatorname{\mathbb{E}}[M_{h+1}|x_{h}]\|^{2}\end{split}$ (61)
@@
=$\overset{(a)}{=}\Pi\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]\Pi(x_{h}-x^{*})+\Pi\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]x^{*}-\Pi\operatorname{\mathbb{E}}[(\widehat{q}_{h}-q)|x_{h}]$
@@
$\operatorname{\mathbb{E}}\|\Pi\operatorname{\mathbb{E}}[M_{h+1}|x_{h}]\|^{2}$ $\leq 3\operatorname{\mathbb{E}}[\|\Pi\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]\Pi(x_{h}-x^{*})\|^{2}]+3\operatorname{\mathbb{E}}[\|\Pi\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]x^{*}\|^{2}]$
$\quad+3\operatorname{\mathbb{E}}[\|\Pi\operatorname{\mathbb{E}}[(\widehat{q}_{h}-q)|x_{h}]\|^{2}]$
$\leq 3\operatorname{\mathbb{E}}[\|\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]\|^{2}\|\Pi(x_{h}-x^{*})\|^{2}]+3\operatorname{\mathbb{E}}[\|\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]\|^{2}\|x^{*}\|^{2}]$
$\quad+3\operatorname{\mathbb{E}}[\|\operatorname{\mathbb{E}}[(\widehat{q}_{h}-q)|x_{h}]\|^{2}]$
@@
Suppose $212\delta_{P}^{2}\leq\lambda_{P}^{2}/$ 2 and $\bar{\beta}$ is small enough such that $\bar{\beta}\leq 1$ and $\bar{\beta}\leq\frac{\lambda_{P}}{2\left((3\sigma_{P}^{2}+2\Lambda_{P}-\lambda_{P}^{2})+\Lambda_{P}\lambda_{P}\right)}$. Then, the above bound gives us
@@ E.2 Proof of Theorem 4(b)
Since $\Pi(\mathbb{E}[x_{h}]-x^{*})\in\mathrm{Ker}(P)_{\perp}$, we can use $A_{6}$ and $A_{8}$ to obtain
@@ F Proof of Theorem 2
Let Assumption 1 hold, and suppose $\Phi$ is full rank. Then $\mathrm{Ker}(M_{\theta})=Z_{\theta}$.
@@
For a large enough $c_{\beta}$, Assumption 2 implies that $\xi^{\top}A_{v}(\theta)\xi\geq(\lambda/2)\|\xi\|^{2}$ for all $\xi=[\eta,\zeta]^{\top}$ such that $\zeta\in\mathrm{Ker}(M_{\theta})_{\perp}$, for all $\theta\in\Theta$.
@@
Recall that $A_{v}(\theta)=\operatorname{\mathbb{E}}_{\theta}[A_{v}(z)]$ where $\operatorname{\mathbb{E}}_{\theta}$ denotes expectation over the distribution of $z=(s,a,s^{\prime})$ where $(s,a)\sim\nu^{\pi_{\theta}}$, $s^{\prime}\sim P(\cdot|s,a)$. Hence, for any $\xi=[\eta,\zeta]$ such that $\zeta\in\mathrm{Ker}(M_{\theta})_{\perp}$, we have
@@
Let $s\in\mathcal{S}_{R}^{\theta}$ and $s^{\prime}\sim P^{\pi_{\theta}}(\cdot\mid s)$ is the next state under the transition kernel induced by policy $\pi_{\theta}$. Then $\mathrm{Ker}(A_{v}(\theta))\subset\mathrm{Ker}(A_{v}(\theta,z))$, where $z=(s,a,s^{\prime})$.
@@
Let $z\in\mathrm{Ker}(A_{v}(\theta))$. Then $z=[0,\,\zeta^{\top}]^{\top}$, where $\zeta\in\mathrm{Ker}(M_{\theta})=Z_{\theta}$. It follows that
@@
G Proof of Theorem 3
We note that condition $(A_{9})$ holds trivially for the NPG update, since $\mathrm{Ker}(A_{u}(\theta))=\{0\}$. Furthermore, by Assumption 4, condition $(A_{8})$ is readily satisfied with $\lambda_{P}=\mu$. Finally, from Assumption 3, we have
@@ G Proof of Theorem 3
=$\overset{(a)}{=}\underbrace{\operatorname{\mathbb{E}}_{u}\left[\frac{1}{B}\sum_{i=1}^{B}\bigg{\{}\eta_{k}^{*}-\eta_{k}+\langle\phi(s_{i+1})-\phi(s_{i}),\zeta_{k}-\zeta_{k}^{*}\rangle\bigg{\}}\nabla_{\theta}\log_{\pi_{\theta_{k}}}(a_{i}|s_{i})\right]}_{T_{0}}+$
@@ I Proof of Theorem 1
We emphasize that our regret dependence on $d$ and $m$ is not directly comparable to the dimension dependence in the linear MDP literature. Our constants align with those in the general parameterized actor-critic framework, where such dependencies may be implicit but are not systematically studied or considered central to theoretical analysis.
Smoothness of the Objective: We note that our analysis assumes smoothness of the average reward objective, $J$. This is a standard assumption in the Policy Gradient (PG) literature and is widely used in both reinforcement learning and even optimization more broadly, as it enables quantifying improvement in the objective after a gradient step.
In the discounted setting, $J$ is well known to be $L$-smooth, with $L$ typically scaling as $(1-\gamma)^{-1}$ (Lemma B.1, [22]). Nearly all works in the average-reward setting assume smoothness, either explicitly, or indirectly by assuming Lipschitz continuity of the value function or by referencing bounds established in the discounted setting [11, 31, 27, 14, 36, 26, 8].
A recent result establishes smoothness in the average-reward, infinite-horizon setting under ergodicity (Theorem 3, [15]), but this relies critically on the assumption of exponentially fast convergence to the stationary distribution and does not apply to general unichain MDPs. In both discounted and average-reward settings, smoothness analyses typically depend on estimating the policy gradient with negligible bias using short, nearly constant-length trajectories. This approach is not viable in the general unichain case, where such rapid convergence cannot be guaranteed.
An alternative analysis [20] shows smoothness of $J$ in a tabular setting with respect to the policy itself, rather than the parameter $\theta$. This approach depends on the linearity of the policy, reward, and transition kernel $P$ in parameter $\alpha$ (see Equations (13) and (14) in [20]), which does not hold under general policy parameterizations.
Given that policy gradient methods are known to perform well across a wide range of environments, we expect a smoothness property, or an analogous result, to hold in the unichain setting. However, establishing this may require a substantially different analytical approach.

### v2

@@ C.3 Proof of Lemma 3
C.4 Bound on the Hitting time of the Recurrent Class
@@
D Analysis of a General Linear Recursion
@@
$(A_{8})\quad x^{\top}Px\geq\lambda_{P}\|x\|^{2},\quad(A_{9})\quad\ker(P)\subseteq\ker(\widehat{P}_{h})$
for all $x\in\ker(P)^{\perp}$ and $h\geq 1$. Let $x^{*}=P^{\dagger}q$, where $P^{\dagger}$ is the Moore-Penrose pseudoinverse of $P$ and $\Pi$ denote the projection operator onto the space $\ker(P)^{\perp}$, i.e., the orthogonal complement of the kernel (null space) of $P$.
@@ D.1 Proof of Theorem 4(a)
Consider a matrix $A$ such that $x^{\top}Ax\geq\lambda\|x\|^{2}$ for all $x\in(\ker(A))^{\perp}$. Then, we have
@@
Let $\Pi$ denote the projection operator onto the space $\ker(P)^{\perp}$, i.e., the orthogonal complement of the kernel (null space) of $P$. Then the following properties hold:
@@
[…]
@@
For the second property, let $\Pi^{\perp}$ denote the projection operator onto $\ker(P)$. By the orthogonal decomposition of the space, we have
$\Pi+\Pi^{\perp}=I,$
@@
$\widetilde{P}x=\widetilde{P}(\Pi x+\Pi^{\perp}x)=\widetilde{P}\Pi x,$
[…]
@@
[…]
@@
[…]
@@
$\overset{(a)}{=}\Pi\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]\Pi(x_{h}-x^{*})+\Pi\operatorname{\mathbb{E}}[(\widehat{P}_{h}-P)|x_{h}]x^{*}-\Pi\operatorname{\mathbb{E}}[(\widehat{q}_{h}-q)|x_{h}]$
@@
[…]
@@
Suppose $12\delta_{P}^{2}\leq\lambda_{P}^{2}/2$ and $\bar{\beta}$ is small enough such that $\bar{\beta}\leq 1$ and $\bar{\beta}\leq\frac{\lambda_{P}}{2\left((3\sigma_{P}^{2}+2\Lambda_{P}-\lambda_{P}^{2})+\Lambda_{P}\lambda_{P}\right)}$. Then, the above bound gives us
@@ D.2 Proof of Theorem 4(b)
Since $\Pi(\mathbb{E}[x_{h}]-x^{*})\in\ker(P)^{\perp}$, we can use $A_{6}$ and $A_{8}$ to obtain
@@ E Proof of Theorem 2
Let Assumption 1 hold, and suppose $\Phi$ is full rank. Then $\ker(M_{\theta})=Z_{\theta}$.
@@
For a large enough $c_{\beta}$, Assumption 2 implies that $\xi^{\top}A_{v}(\theta)\xi\geq(\lambda/2)\|\xi\|^{2}$ for all $\xi=[\eta,\zeta]^{\top}$ such that $\zeta\in\ker(M_{\theta})^{\perp}$, for all $\theta\in\Theta$.
@@
Recall that $A_{v}(\theta)=\operatorname{\mathbb{E}}_{\theta}[A_{v}(z)]$ where $\operatorname{\mathbb{E}}_{\theta}$ denotes expectation over the distribution of $z=(s,a,s^{\prime})$ where $(s,a)\sim\nu^{\pi_{\theta}}$, $s^{\prime}\sim P(\cdot|s,a)$. Hence, for any $\xi=[\eta,\zeta]$ such that $\zeta\in\ker(M_{\theta})^{\perp}$, we have
@@
Let $s\in\mathcal{S}_{R}^{\theta}$ and $s^{\prime}\sim P^{\pi_{\theta}}(\cdot\mid s)$ is the next state under the transition kernel induced by policy $\pi_{\theta}$. Then $\ker(A_{v}(\theta))\subset\ker(A_{v}(\theta,z))$, where $z=(s,a,s^{\prime})$.
@@
Let $z\in\ker(A_{v}(\theta))$. Then $z=[0,\,\zeta^{\top}]^{\top}$, where $\zeta\in\ker(M_{\theta})=Z_{\theta}$. It follows that
@@
F Proof of Theorem 3
We note that condition $(A_{9})$ holds trivially for the NPG update, since $\ker(A_{u}(\theta))=\{0\}$. Furthermore, by Assumption 4, condition $(A_{8})$ is readily satisfied with $\lambda_{P}=\mu$. Finally, from Assumption 3, we have
@@ F Proof of Theorem 3
$\overset{(a)}{=}\underbrace{\operatorname{\mathbb{E}}_{u}\left[\frac{1}{B}\sum_{i=1}^{B}\bigg\{\eta_{k}^{*}-\eta_{k}+\langle\phi(s_{i+1})-\phi(s_{i}),\zeta_{k}-\zeta_{k}^{*}\rangle\bigg\}\nabla_{\theta}\log_{\pi_{\theta_{k}}}(a_{i}|s_{i})\right]}_{T_{0}}+$
@@ H Proof of Theorem 1
Remark 2 (Scalability of $C_{\mathrm{hit}}$ and $C_{\mathrm{tar}}$).
The constants $C_{\mathrm{hit}}$ and $C_{\mathrm{tar}}$ reflect the difficulty of exploration under a given policy class and may scale with structural properties of the MDP such as state-space size or connectivity. Such dependence is inherent, as any algorithm must visit all states to avoid constant per-iteration regret, and even in ergodic MDPs the mixing time $t_{\mathrm{mix}}$ can scale with the state space. Our result extends existing ergodic-MDP guarantees to the more general unichain setting without degrading these constants. Specifically, while state-of-the-art actor–critic methods achieve complexity $O(t_{\mathrm{mix}}^{3}\sqrt{T})$ [16], our bound $O(C^{3}\sqrt{T})$ involves a constant $C$ that characterizes the Cesàro mixing time [22, Sec. 6.6], which satisfies $C\leq 7t_{\mathrm{mix}}$ for ergodic chains [22, Ex. 6.11]. In certain Markov chains, such as the biased random walk on the $n$-cycle [22, Ex. 24.2], the Cesàro mixing time is only $O(n)$ versus the standard mixing time which is $\Theta(n^{2})$, illustrating that averaging distributions over time accelerates convergence compared to single-step mixing.
Remark 3 (On the Unichain Assumption).
To the best of our knowledge, no existing policy gradient methods (with general policy parametrization) have theoretical guarantees for average-reward, infinite-horizon MDPs beyond the unichain case. Our work is the first to establish such guarantees under the general unichain assumption, which is the weakest known condition under which the policy gradient theorem holds [33]. While value-based methods admit guarantees beyond this setting, policy gradient approaches require fundamentally different analyses, making existing techniques inapplicable. A possible, though still conjectural, direction for relaxing the unichain assumption is to restrict policy search to unichain-inducing policies by maintaining strong exploration initially and annealing it as the algorithm converges. Notably, weakly communicating MDPs, the most general class solvable from a single stream of experience, always admit an optimal unichain policy [30].